Calculate the pH of 50 ml of 0.1000 M methylamine solution after the following volume of 0.1000 M HNO3, where added 60

1
Expert's answer
2022-06-16T01:32:15-0400

n(CH3NH2) = CM * V = 0.1 * 0.05 = 0.005mol

n(HNO3) = CM * V = 0.1 * 0.06 = 0.006mol


CH3NH2 + HNO3 = CH3OH + N2 + H2O

1mol — 1mol

0.005mol — X


X = 0.005 mol


n(HNO3) = 0.006 - 0.005 = 0.001 mol

Vnew = 0.05 + 0.06 = 0.11 L


CM(HNO3) =n/V=0.001/0.11 =0.0091M


HNO3 => H+ + NO3-

1mol HNO3 => 1mol H+

0.0091 M => X = 0.0091 M H+


pH = -log [H+] = - log (9.1*10-3) = 2.04


Answer: pH = 2.04